At the time of the static budget, Burlington Burrito expected to spend $30,550 on labor. In that budget, they expected that each burrito meal would require an average of 10 minutes to make and that each hour of labor would cost $6.50.

During the period, they spent $33,000 to make 30,000 burrito meals, and employees worked a total of 4,600 hours.

Explain the $2,450 difference between the static budget and actual cost, using the three variances (Rate, Usage and Volume Variances). Make sure you include if they are favorable or unfavorable variances.
Answer #1

Given data

a. Standard labor cost for 30,000 burrito meals = $30,550

b. Standard rate per labor hour = $6.50

c. Therefore standard hours required to produce 30,000 burrito meals (a/b) = $30,550 / $6.50 = 4700 hours

d. Actual labor cost for 30,000 burrito meals = $33,000

e. Actual labor hours worked to produce 30,000 burrito meals = 4,600 hours

f. Therefore actual rate per labor hour (d/e) = $33,000 / 4,600 hours = $7.1739

1. Labor Rate Variance (LRV) = (Standard rate per labor hour - Actual rate per labor hour) * Actual labor hours worked

= ($6.50 - $7.1739) * 4,600 hours = $3,100 (U) U represents Unfavorable Variance

2. Labor Efficiency Variance (LEV) = (Standard labor hours - Actual labor hours) * Standard rate per labor hour

= (4,700 hours - 4,600 hours) * $6.50 = $650 (F) F represents Favorable Variance

Therefore, from the above it can be seen that the difference of $2,450 (U) ($3,100 - $650), arising because of the labor rate per hour for the work they were engaged. The labor has been efficiently worked during the production which results in better results.